# Survey on the Evaluation of Generative Models in Music

### v3

@@ 1. Introduction
In recent years, advances in system architecture and training methodologies of generative systems in machine learning have led to increasingly powerful systems that have been applied to a variety of tasks and use cases. Particularly prominent are systems in natural language processing (e.g., (Brown et al., 2020)) and image models (e.g., (Karras et al., 2019; Ramesh et al., 2021)). Generative systems for music, while not featured as prominently in the media, have also seen considerable progress. Although the concept of computer-generated music dates back to the mid-20th century (Hiller and Isaacson, 1959; Supper, 2001) and has been actively researched since then (Cope, 2000; Fernandez and Vico, 2013), there has been a rapid increase in interest and publications in the past five years (Civit et al., 2022), mostly driven by advances in neural approaches. Inspired by Herremans et al. (2017), we categorize generative music systems based on (i) architecture, (ii) output, and (iii) input or control. [^1] References will be limited to a select number of representative systems. A meta-review by Civit et al. (2022) lists the prevalent architectures with decreasing number of occurrences as Recurrent Neural Networks (RNNs) (Hadjeres et al., 2017; Mao et al., 2018), Feed-Forward Networks ($F$) (Yang et al., 2017; Dong et al., 2018), Variational Audio Encoders (VAEs) (Roberts et al., 2018; Dhariwal et al., 2020), evolutionary algorithms (Stoltz and Aravind, 2019; De Prisco et al., 2020), and Transformer-based approaches (Huang et al., 2019; Donahue et al., 2019; Copet et al., 2023), sometimes combined with Generative Adversarial Networks (GANs) (Yang et al., 2017; Dong et al., 2018), with Diffusion networks (Evans et al., 2024, 2025; Liu et al., 2023; Chen et al., 2024; Schneider et al., 2024; Ning et al., 2025), or with Rectified Flows (Li et al., 2025). Details of many typical neural architectures for music generation can be found in Briot et al. (2020).
@@
Figure 1. Structure of the presented evaluation approaches and methodologies.
[…]
Structural flow-chart overview of the topics covered in the paper, starting with evaluation targets (musical qualities, audio fidelity, originality, and conditioning for output evaluation and ease of use, creative experience, user impression, and musical practice for evaluation of usability and user experience. Subsequent steps outline the detailed methodologies usable for assessment.
@@ 2.1.4. Conditioning
$2.$ Usability & user experience
@@ 3.2.1. Methodology
Figure 2. Overview of the evaluation methodologies for generative music systems, including both reference-dependent and reference-independent approaches.
[…]
The evaluation methodologies structured into a coherent framework that integrates different approaches.
@@ 4.2. User data collection: Quantitative
User experience-oriented scales include the User Experience Questionnaire (Laugwitz et al., 2008) and the User Engagement Scale (O’Brien et al., 2018). The most influential questionnaire with respect to creative technology within the last 10 years is arguably the Creativity Support Index (CSI) (Cherry and Latulipe, 2014): a questionnaire designed to test the capacity of a tool to support creativity, offering factors for several important aspects of the creative user experience including: focused attention, enjoyment and collaboration. For AI music, however, we found surprisingly few examples of AI music user studies that have adopted the CSI — Bougueng Tchemeube et al. (2023) a notable exception. We instead observe many examples where researchers have chosen to define their own questionnaires to explore constructs for the AI systems that are not captured in current standardized scales. For instance, Louie et al. (2020) invented their own questions to capture a person’s feelings of agency. Ford and Bryan-Kinns (2023) identified reflection as an aspect missing from the CSI whilst being an important factor in AI music making (Ford et al., 2024; Ford and Bryan-Kinns, 2022). Alongside established measures of creativity support and usability, Bougueng Tchemeube et al. (2023) also added questions central to human-AI interaction on trust, perceived authorship, and flexibility. This use of researcher defined questionnaires reflects the dominance of measures of engagement and usability in assessing AI interaction in creativity-related HCI research (Cox et al., 2025; Rhys Cox et al., 2025).
@@ 4.3.3. First-person perspectives
$4.$ Other HCI evaluation approaches
@@ 5.2. Perceptual and musical relevance of objective metrics
Perceptual studies of specific descriptors are necessary to understand their relevance and meaning. Simply finding a difference between two set of data with respect to one descriptor does not automatically mean that these data are different from a perceptual point of view. Recent studies on the suitability of learned embeddings for evaluation seem to exclusively focus on overall relevance for summary aesthetic judgments without considering interpretability or musical meaning (Huang et al., 2025; Grötschla et al., 2025).
@@ 6. Conclusion
We presented an overview of the state-of-the-art in evaluating generative systems in music from the perspective of both the system output and the usability of the system. We categorized different evaluation goals and targets, as well as corresponding methodologies and metrics, and concluded that the current state of system assessment makes it difficult to generalize results, compare state of the art systems objectively, and measure progress in the field. The main challenges identified include (i) the perceptual and musical meaningfulness of current evaluation metrics, (ii) the internal and external validity of common experimental setups, and (iii) the lack of reproducibility of experiments, emphasizing the need for de-facto evaluation standards adopted by the research community. While the direct assessment of aesthetics might “remain out of reach in our lifetime and perhaps forever” (Galanter, 2012), there is a need for methodologies and metrics that give us at least a glimpse into some aspects of quality.
In addition to the evaluation targets presented above, there exist other evaluation targets of interest. Of note could be targets that can be summarized under the umbrella of Responsible AI (Piskopani et al., 2023). These include, e.g., (i) explainability: the increasing complexity of machine learning systems puts forth questions with respect to usage and deployment of these systems (Kambhampati, 2022), including in the arts (Bryan-Kinns, 2024), however, standardized evaluation strategies do not exist (Batlle-Roca et al., 2023); (ii) bias: bias of machine learning systems is a known problem (Barnett, 2023; Bryan-Kinns et al., 2024c) and can —in the case of generative music systems— lead to marginalization of non-mainstream musical styles (Bryan-Kinns et al., 2024a); (iii) ethical use of data: ethical data acquisition approaches, guiding principles, and transparency on data content and curation are crucial for the holistic evaluation of a machine learning system (Jo and Gebru, 2020), as the call for “dataset audit cards” (Birhane and Prabhu, 2021) and formalized datasheets for datasets (Srinivasan et al., 2021; Gebru et al., 2021) emphasize; (iv) resource use(Strubell et al., 2019; Chien et al., 2023; Berthelot et al., 2024): the high energy consumption of today’s models can be linked to non-negligible environmental impacts, and the reporting of carbon emissions (Henderson et al., 2020) or the relation of energy consumption to the subjective output quality of generative music systems (Douwes et al., 2023) could promote energy-responsible research.

### v4

@@ 1. Introduction
In recent years, advances in system architecture and training methodologies of generative systems in machine learning have led to increasingly powerful systems that have been applied to a variety of tasks and use cases. Particularly prominent are systems in natural language processing (e.g., (Brown et al., 2020)) and image models (e.g., (Karras et al., 2019; Ramesh et al., 2021)). Generative systems for music, while not featured as prominently in the media, have also seen considerable progress. Although the concept of computer-generated music dates back to the mid-20th century (Hiller and Isaacson, 1959; Supper, 2001) and has been actively researched since then (Cope, 2000; Fernandez and Vico, 2013), there has been a rapid increase in interest and publications in the past five years (Civit et al., 2022), mostly driven by advances in neural approaches. Inspired by Herremans et al. (2017), we categorize generative music systems based on (i) architecture, (ii) output, and (iii) input or control. [^1] References will be limited to a select number of representative systems. A meta-review by Civit et al. (2022) lists the prevalent architectures with decreasing number of occurrences as Recurrent Neural Networks (RNNs) (Hadjeres et al., 2017; Mao et al., 2018), Feed-Forward Networks (FF) (Yang et al., 2017; Dong et al., 2018), Variational Audio Encoders (VAEs) (Roberts et al., 2018; Dhariwal et al., 2020), evolutionary algorithms (Stoltz and Aravind, 2019; De Prisco et al., 2020), and Transformer-based approaches (Huang et al., 2019; Donahue et al., 2019; Copet et al., 2023), sometimes combined with Generative Adversarial Networks (GANs) (Yang et al., 2017; Dong et al., 2018), with Diffusion networks (Evans et al., 2024, 2025; Liu et al., 2023; Chen et al., 2024; Schneider et al., 2024; Ning et al., 2025), or with Rectified Flows (Li et al., 2025). Details of many typical neural architectures for music generation can be found in Briot et al. (2020).
@@
Figure 1. Structure of the presented evaluation approaches and methodologies.:
@@ 2.1.4. Conditioning
2.2. Usability & user experience
@@ 3.2.1. Methodology
Figure 2. Overview of the evaluation methodologies for generative music systems, including both reference-dependent and reference-independent approaches.:
@@ 4.2. User data collection: Quantitative
User experience-oriented scales include the User Experience Questionnaire (Laugwitz et al., 2008) and the User Engagement Scale (O’Brien et al., 2018). The most influential questionnaire with respect to creative technology within the last 10 years is arguably the Creativity Support Index (CSI) (Cherry and Latulipe, 2014): a questionnaire designed to test the capacity of a tool to support creativity, offering factors for several important aspects of the creative user experience including: focused attention, enjoyment and collaboration. For AI music, however, we found surprisingly few examples of AI music user studies that have adopted the CSI — Bougueng Tchemeube et al. (2023) a notable exception. We instead observe many examples where researchers have chosen to define their own questionnaires to explore constructs for the AI systems that are not captured in current standardized scales. For instance, Louie et al. (2020) invented their own questions to capture a person’s feelings of agency. Ford and Bryan-Kinns (2023) identified reflection as an aspect missing from the CSI whilst being an important factor in AI music making (Ford et al., 2024; Ford and Bryan-Kinns, 2022). Alongside established measures of creativity support and usability, Bougueng Tchemeube et al. (2023) also added questions central to human-AI interaction on trust, perceived authorship, and flexibility. This use of researcher defined questionnaires reflects the dominance of measures of engagement and usability in assessing AI interaction in creativity-related HCI research (Rhys Cox et al., 2025b, a).
@@ 4.3.3. First-person perspectives
4.4. Other HCI evaluation approaches
@@ 5.2. Perceptual and musical relevance of objective metrics
Perceptual studies of specific descriptors are necessary to understand their relevance and meaning. Simply finding a difference between two set of data with respect to one descriptor does not automatically mean that these data are different from a perceptual point of view. Recent studies on the suitability of learned embeddings for evaluation seem to focus on overall relevance for summary aesthetic judgments without considering interpretability or musical meaning (Huang et al., 2025; Grötschla et al., 2025).
@@ 6. Conclusion
We presented an overview of the state-of-the-art in evaluating generative systems in music from the perspective of both the system output and the usability of the system. We categorized different evaluation goals and targets, as well as corresponding methodologies and metrics, and concluded that the current state of system assessment makes it difficult to generalize results, compare state of the art systems objectively, and measure progress in the field. The main challenges identified include (i) the perceptual and musical meaningfulness of current evaluation metrics, (ii) the internal and external validity of common experimental setups, and (iii) the lack of reproducibility, emphasizing the need for de-facto evaluation standards adopted by the research community. While the direct assessment of aesthetics might “remain out of reach in our lifetime and perhaps forever” (Galanter, 2012), there is a need for methodologies and metrics that give us at least a glimpse into some aspects of quality.
In addition to the evaluation targets presented above, there exist other evaluation targets of interest. Of note could be targets that can be summarized under the umbrella of Responsible AI (Piskopani et al., 2023). These include, e.g., (i) explainability: the increasing complexity of machine learning systems puts forth questions with respect to usage and deployment of these systems (Kambhampati, 2022), including in the arts (Bryan-Kinns, 2024), however, standardized evaluation strategies do not exist (Batlle-Roca et al., 2023); (ii) bias: bias of machine learning systems is a known problem (Barnett, 2023; Bryan-Kinns et al., 2024c) and can —in the case of generative music systems— lead to marginalization of non-mainstream musical styles (Bryan-Kinns et al., 2024a); (iii) ethical use of data: ethical data acquisition, guiding principles, and transparency on data content and curation are crucial for the holistic evaluation of a machine learning system (Jo and Gebru, 2020), as the call for “dataset audit cards” (Birhane and Prabhu, 2021) and formalized datasheets for datasets (Srinivasan et al., 2021; Gebru et al., 2021) emphasize; (iv) resource use(Strubell et al., 2019; Chien et al., 2023; Berthelot et al., 2024): the high energy consumption of today’s models can be linked to environmental impacts; the reporting of carbon emissions (Henderson et al., 2020) or the relation of energy consumption to the subjective output quality of generative music systems (Douwes et al., 2023) could promote energy-responsible research.